### v3

@@ Abstract
  • The subclass of $\mathsf{VP}$ that contains polynomial families with bounded coefficients, has efficient equations. Over finite fields, this result holds without any restriction on coefficients. Further, both these results also extend to the class $\mathsf{VNP}$ as is.
  • Over fields of characteristic zero, $\mathsf{VNP}$ does not have any efficient equations, if the Permanent is exponentially hard for algebraic circuits.
@@
The first set of results follows essentially by algebraizing the well-known method of generating hardness from non-trivial hitting sets (e.g. Heintz and Schnorr 1980). The conditional hardness of equations for $\mathsf{VNP}$ uses the fact that pseudorandomness against a class can be extracted from a polynomial that is (sufficiently) hard for that class (Kabanets and Impagliazzo, 2004).
[…]
@@ 1.1 The Natural Proofs framework of Razborov and Rudich
A proof that a certain family of Boolean functions cannot be computed by circuits in $\mathscr{C}$ is said to be a natural lower bound proof if it (perhaps implicitly) proceeds via establishing a natural property useful against $\mathscr{C}$, and showing that the candidate hard function has this property. Razborov and Rudich then showed that most of the Boolean circuit lower bound proofs that we know (for example, lower bounds for $\mathsf{AC}^{0}$ circuits [FSS84, H86] or lower bounds for $\mathsf{AC}^{0}[\oplus]$ circuits [R87, S87]) fit into this framework, maybe with some work, and hence are natural in this sense. Further they argue that, under standard cryptographic assumptions, the proof of a lower bound against any sufficiently rich circuit class (such as the class $\mathsf{P}/\operatorname{poly}$) cannot be natural! Thus, under standard cryptographic assumptions, most of the current lower bound techniques are not strong enough to show super-polynomial lower bounds for general Boolean circuits.
@@ 1.2 Algebraically Natural Proofs
Algebraic complexity is the study of computational questions about polynomials as formal objects. The basic model of computation here, an algebraic circuit, is an algebraic analogue of a boolean circuit with the gates of the circuit being labeled by $+$ (sum) and $\times$ (product) gates as opposed to Boolean functions; the size of a circuit is the number of wires (edges) in it. [^3] See 2.5 for a formal definition. The algebraic analogue of $\mathsf{P}/\operatorname{poly}$ is the class $\mathsf{VP}$ of polynomial families $\{f_{n}\}$, where $f_{n}$ is an $n$ variate polynomial of degree and algebraic circuit size $\operatorname{poly}(n)$. A fundamental question in this setting is to come up with explicit families of polynomials, those in the class $\mathsf{VNP}$ (the algebraic analog of $\mathsf{NP}/\operatorname{poly}$), which are not in $\mathsf{VP}$. While the state of the art of size lower bounds for algebraic circuits is a bit better than that for Boolean circuits, with slightly super linear lower bounds having been shown by Strassen [S73] and Baur & Strassen [BS83], this lower bound has seen no improvements for nearly four decades. The recent breakthrough by Limaye, Srinivasan and Tavenas [LST21] does indeed suggest that stronger lower bounds might be within reach in the near future, but it is far from clear how that could be done. This absence of progress has led to some research towards understanding the limitations of the current proof techniques in proving strong lower bounds for algebraic circuits.
@@
In the last few years, this question of trying to find an algebraic analogue of the barrier results in [RR97] has received substantial attention. It was observed by various authors [AD08, G15, FSV18, GKSS17] that most of the currently known proofs of algebraic circuit lower bounds fit into a common unifying framework, not unlike that in [RR97], although of a more algebraic nature. Indeed, these proofs also implicitly go via defining a property for the set of all polynomials and using this property to separate the hard polynomial from the easy ones. Moreover, the notions of largeness and constructibility in 1.1 also seem to extend to these proofs.
@@
For some $n,d\in\mathbb{N}$, let $\pazocal{C}_{n,d}$ be a set of $n$-variate polynomials of total degree at most $d$; i.e. $\pazocal{C}_{n,d}\subseteq\mathbb{F}[\mathbf{x}]^{\leq d}$. Then, for $N=\binom{n+d}{n}$, a nonzero polynomial $P_{N}(\mathbf{Z})$ is said to be an equation for $\pazocal{C}_{n,d}$ if for all $f(\mathbf{x})\in\pazocal{C}_{n,d}$, we have that $P_{N}(\overline{\operatorname{coeff}}(f))=0$, where $\overline{\operatorname{coeff}}(f)$ is the coefficient vector of $f$.
@@
Analogous to the abstraction of natural proofs for Boolean circuit lower bounds, this framework of algebraically natural proofs turns out to be rich and general enough that almost all of our current proofs of algebraic circuit lower bounds are in fact algebraically natural, or can be viewed in this framework with a little work [G15]. Thus, this definition seems like an important first step towards understanding the strengths and limitations of many of our current lower bound techniques in algebraic complexity.
@@ 1.4 Discussion and relations to prior work
Bläser, Ikenmeyer, Jindal and Lysikov [BIJL18] studied the complexity of equations in a slightly different context. They draw connections between the existence of efficiently constructible equations of a variety and the problem of testing (non)membership in it and use the conditional hardness of the (non)membership testing problem for certain varieties to rule out the existence of efficiently computable equations for them. More precisely, they show that if all the equations for the variety of matrices with zero permanent are constructible by small constant-free algebraic circuits, then the non-membership problem for this variety can be decided in the class $\exists\mathsf{BPP}$. Thus, unless $\mathsf{P}^{\#\mathsf{P}}\subseteq\exists\mathsf{BPP}$, the equations of this variety do not have small, low degree constant free algebraic circuits. In a subsequent work ([BILPS19]), the results of [BIJL18] are generalized to min-rank or slice-rank varieties. However, in the bounded coefficient setting (and over finite fields), our results show that the contrary is true, and $\mathsf{VP}$ does have efficiently computable low degree equations. We also note that the set-up in these papers differ from that in our paper, and that of [GKSS17, FSV18]. One way to interpret this difference is that [BIJL18] shows that “variety of small completion rank tensors” cannot be “cut out” by efficient equations, whereas [GKSS17, FSV18] and our paper asks if every equation for this variety requires large complexity.
@@
For Boolean circuits, Chow [C11] shows a way of circumventing the natural proofs’ barrier in [RR97] by providing (under standard cryptographic assumptions) an explicit almost natural proof that is useful against $P/\operatorname{poly}$ as well as constructive in nearly linear time, but compromises on the largeness condition. Furthermore, Chow [C11] shows the unconditional existence of a natural property useful against $P/\operatorname{poly}$ (infinitely often) constructive in linear size that has a weakened largeness condition. In some sense, Theorem 1.7 and Theorem 1.6 are analogous to the work of Chow [C11], albeit in the algebraic world.
@@ 1.5 Proof ideas
In order to show that the equations are non-trivial, in the sense that there exist polynomials with bounded integer coefficients which do not pass this test, we need to show that there are nonzero polynomials with bounded integer coefficients which vanish everywhere on the hitting set $\pazocal{H}$. We show this via a well-known lemma of Siegel [^6] A statement of the lemma can be found here. Refer to [S14a] for details., which uses a simple pigeon-hole argument to show that an under-determined system of homogeneous linear equations where the constraint matrix has small integer entries has a nonzero solution with small integer entries.
@@
As was observed in [FSV18, GKSS17], a lower bound for equations for a class of polynomials is equivalent to showing the existence of succinctly describable hitting sets for this class. For our proof of LABEL:thm:lb-on-equations-for-vnp-slice we show that, assuming that the permanent is sufficiently hard, the coefficient vectors of polynomials in $\mathsf{VNP}$ form a hitting set for the class $\mathsf{VP}$. The connection between hardness and randomness in algebraic complexity is well known via a result of Kabanets and Impagliazzo [KI04], and we use this connection for our proof, along with some additional ideas. It is useful to note that the above-mentioned result of Kabanets and Impagliazzo [KI04] is essentially an algebraic analogue of the hardness vs randomness paradigm introduced by Nisan and Wigderson [NW94] in the boolean world. We briefly describe a high level sketch of our proof in a bit more detail now.
@@
For the proof of LABEL:thm:lb-on-equations-for-vnp-slice, we show that the points in the image of the map $\mathsf{Gen}_{\operatorname{Perm}}$, can be viewed as the coefficient vectors of polynomials in $\mathsf{VNP}$, or, equivalently in the terminology in [FSV18, GKSS17], that the Kabanets-Impagliazzo hitting set generator is $\mathsf{VNP}$-succinct. To this end, we work with a specific instantiation of the construction of the Kabanets-Impagliazzo generator where the underlying construction of combinatorial designs is based on Reed-Solomon codes. Although this is perhaps the most well known construction of combinatorial designs, there are other (and in some parameters, better) constructions known. However, our proof relies on the properties of this particular construction to obtain the succinct description. Our final proof is fairly short and elementary, and is based on extremely simple algebraic ideas and making generous use of the fact that we are trying to prove a lower bound for equations for $\mathsf{VNP}$ and not $\mathsf{VP}$.
@@
We begin with some notations and preliminaries in Section 2. In Section 3, we outline the existence of efficient hitting sets, over characteristic-zero fields, for any class that has a low-degree, low-variate universal polynomials; $\mathsf{VP}$ and $\mathsf{VNP}$ are examples of such classes. Existence of efficient hitting sets for $\mathsf{VP}$ and $\mathsf{VNP}$ over finite fields is outlined in Section 3.2. Note that all the arguments in Section 3 are present in previous works, but the results over characteristic zero have not been stated in this generality prior to this work, to the best of our knowledge. In Section 4, we use the above results to first prove Theorems 1.6 to 1.9. We then show the conditional hardness of equations for $\mathsf{VNP}$ in Section 5. Finally, we provide some open questions that arise from our work, and the prior literature around algebraic natural proofs in LABEL:sec:open-questions.
@@ 2.1 Notation and basics
Lemma 2.3 (Polynomial Identity Lemma [O22, DL78, S80, Z79]).
@@ 2.2.2 Equations, natural proofs and succinct hitting sets over characteristic zero
$\mathsf{VP}$ has $\mathsf{VP}$-natural proofs normal-$\Longleftrightarrow$ $\mathsf{VP}$ does not have $\mathsf{VP}$-succinct hitting sets.
@@ 2.3 Universal Circuits
Let $f$ be an $n$-variate degree $d$ polynomial computable by a circuit $C$ of size $s$. Using the classical depth reduction result due to Valiant et al. [VSBR83], $f$ has a circuit $C^{\prime}$ of size $s^{\prime}=\operatorname{poly}(n,d,s)$ and depth $\ell=O(\log d)$ with the following properties (see, e.g., [S15] for a complete proof).
@@ 2.4 Hardness-Randomness Connections
Kabanets and Impagliazzo [KI04] obtain hitting set generators from polynomials that are hard to compute for algebraic circuits. The following lemma is crucial to the proof of LABEL:thm:lb-on-equations-for-vnp-slice.
@@ 3.2 Finite fields: Hitting Sets for VP and VNP
Lemma 3.8 (Folklore (cf. Forbes [F14, Lemma 3.2.14])).
@@ 3.3 Finite fields: Hitting Sets for VNP
In order to prove the existence of a hitting set for the class $\pazocal{D}(n,d,s)$, we will need a bound on the number of polynomials in the class $\pazocal{D}(n,d,s)$ as well as a bound on the size of an explicit hitting set for the class of $n$-variate degree at most $d$ polynomials. These two bounds are summarized in the following claims, proofs of which can be found in [F14].
Claim 3.10 (Lemma 3.1.6 in [F14]).
@@
The rest of the proof follows exactly along the lines of the proof of Lemma 3.2.14 in [F14].

### v4

@@ Abstract
  • The subclass of $\mathsf{VP}$ that contains polynomial families with bounded coefficients, has efficient equations. Over finite fields, this result holds without any restriction on coefficients. Further, both these results extend to any class that admits a low-variate, low-degree universal map: a generator for all polynomials in the class. Most well-studied classes have this property, e.g. $\mathsf{VNP}$, $\mathsf{VBP}$, $\mathsf{VF}$.
  • Over fields of characteristic zero, $\mathsf{VNP}$ does not have any efficient equations, if the Permanent is exponentially hard for algebraic circuits. Moreover, exponential hardness of the Permanent in the approximative sense, even rules out efficient equations of large degree.
@@
The first set of results is obtained by algebraizing the well-known method of generating hardness from non-trivial hitting sets, and by generalizing the result of Heintz and Schnorr (1980) that proves the existence of hitting sets for $\mathsf{VP}$. The conditional hardness of equations for $\mathsf{VNP}$ uses the fact that pseudorandomness against a class can be extracted from a polynomial that is (sufficiently) hard for that class (Kabanets and Impagliazzo, 2004).
@@ 1.1 The Natural Proofs framework of Razborov and Rudich
A proof that a certain family of Boolean functions cannot be computed by circuits in $\mathscr{C}$ is said to be a natural lower bound proof if it (perhaps implicitly) proceeds via establishing a natural property useful against $\mathscr{C}$, and showing that the candidate hard function has this property. Razborov and Rudich then showed that most of the Boolean circuit lower bound proofs that we know (for example, lower bounds for $\mathsf{AC}^{0}$ circuits [FSS84, Hås86] or lower bounds for $\mathsf{AC}^{0}[\oplus]$ circuits [Raz87, Smo87]) fit into this framework, maybe with some work, and hence are natural in this sense. Further they argue that, under standard cryptographic assumptions, the proof of a lower bound against any sufficiently rich circuit class (such as the class $\mathsf{P}/\operatorname{poly}$) cannot be natural! Thus, under standard cryptographic assumptions, most of the current lower bound techniques are not strong enough to show super-polynomial lower bounds for general Boolean circuits.
@@ 1.2 Algebraically Natural Proofs
Algebraic complexity is the study of computational questions about polynomials as formal objects. The basic model of computation here, an algebraic circuit, is an algebraic analogue of a boolean circuit with the gates of the circuit being labeled by $+$ (sum) and $\times$ (product) gates as opposed to Boolean functions; the size of a circuit is the number of wires (edges) in it. [^3] See 2.5 for a formal definition. The algebraic analogue of $\mathsf{P}/\operatorname{poly}$ is the class $\mathsf{VP}$ of polynomial families $\{f_{n}\}$, where $f_{n}$ is an $n$ variate polynomial of degree and algebraic circuit size $\operatorname{poly}(n)$. A fundamental question in this setting is to come up with explicit families of polynomials, those in the class $\mathsf{VNP}$ (the algebraic analog of $\mathsf{NP}/\operatorname{poly}$), which are not in $\mathsf{VP}$. While the state of the art of size lower bounds for algebraic circuits is a bit better than that for Boolean circuits, with slightly super linear lower bounds having been shown by Strassen [Str73] and Baur & Strassen [BS83], this lower bound has seen no improvements for nearly four decades. The recent breakthrough by Limaye, Srinivasan and Tavenas [LST24] does indeed suggest that stronger lower bounds might be within reach in the near future, but it is far from clear how that could be done. This absence of progress has led to some research towards understanding the limitations of the current proof techniques in proving strong lower bounds for algebraic circuits.
@@
In the last few years, this question of trying to find an algebraic analogue of the barrier results in [RR97] has received substantial attention. It was observed by various authors [AD08, Gro15, FSV18, GKSS17] that most of the currently known proofs of algebraic circuit lower bounds fit into a common unifying framework, not unlike that in [RR97], although of a more algebraic nature. Indeed, these proofs also implicitly go via defining a property for the set of all polynomials and using this property to separate the hard polynomial from the easy ones. Moreover, the notions of largeness and constructibility in 1.1 also seem to extend to these proofs.
@@
For some $n,d\in\mathbb{N}$, let $\pazocal{C}_{n,d}$ be a set of $n$-variate polynomials of total degree at most $d$; i.e. $\pazocal{C}_{n,d}\subseteq\mathbb{F}[\mathbf{x}]^{\leq d}$.
Then, for $N=\binom{n+d}{n}$, a nonzero polynomial $P_{N}(\mathbf{Z})$ is said to be an equation for $\pazocal{C}_{n,d}$ if for all $f(\mathbf{x})\in\pazocal{C}_{n,d}$, we have that $P_{N}(\overline{\operatorname{coeff}}(f))=0$, where $\overline{\operatorname{coeff}}(f)$ is the coefficient vector of $f$.
@@
Analogous to the abstraction of natural proofs for Boolean circuit lower bounds, this framework of algebraically natural proofs turns out to be rich and general enough that almost all of our current proofs of algebraic circuit lower bounds are in fact algebraically natural, or can be viewed in this framework with a little work [Gro15]. Thus, this definition seems like an important first step towards understanding the strengths and limitations of many of our current lower bound techniques in algebraic complexity.
@@ 1.3 Our results
Conditional approximative hardness of equations for VNP
Finally, we show that the statement can be strengthened further if we assume that ${\operatorname{Perm}_{n}}$ is hard even in the approximative sense.
Definition 1.12 (Approximative circuits and complexity).
A circuit $C(\mathbf{x},\varepsilon)$ is said to approximately compute an $n$-variate polynomial $f(\mathbf{x})$ if it satisfies $C(\mathbf{x},\varepsilon)=\varepsilon^{M}\cdot f(\mathbf{x})+\varepsilon^{M+1}\cdot g(\mathbf{x},\varepsilon)\in\mathbb{F}(\varepsilon)[\mathbf{x}]$ for some $(n+1)$-variate polynomial $g$, and $M\in\mathbb{N}$.
For a polynomial $f(\mathbf{x})$, its approximative complexity is defined to be the size of the smallest circuit that approximately computes it.
It is clear that the approximative complexity of $f$ is at most $\operatorname{CircuitSize}(f)$, just because one can take $M$ to be 00, and $g(\mathbf{x},\varepsilon)$ to be the identically zero polynomial in 1.12.
We show that if we assume that ${\operatorname{Perm}_{n}}$ is exponentially hard even in the approximative sense, then we can rule out all efficient equations for $\mathsf{VNP}$, irrespective of their degree. More formally, we show the following.
Theorem 1.13 (Conditional Approximative Hardness of Equations for VNP).
Let $\varepsilon>0$ be a constant. Suppose that the permanent family $\operatorname{Perm}_{m}$ requires approximative circuits of size $2^{m^{\varepsilon}}$.
Then $\mathsf{VP}_{\mathrm{nb}}$ has $\mathsf{VNP}$-succinct hitting sets, and so there are no efficiently computable equations for $\mathsf{VNP}$.
This is an important statement, especially when we think of degree as a measure of the “largeness” of the purported equation.
@@ 1.4 Discussion and relations to prior work
Bläser, Ikenmeyer, Jindal and Lysikov [BIJL18] studied the complexity of equations in a slightly different context. They draw connections between the existence of efficiently constructible equations of a variety and the problem of testing (non)membership in it and use the conditional hardness of the (non)membership testing problem for certain varieties to rule out the existence of efficiently computable equations for them. More precisely, they show that if all the equations for the variety of matrices with zero permanent are constructible by small constant-free algebraic circuits, then the non-membership problem for this variety can be decided in the class $\exists\mathsf{BPP}$. Thus, unless $\mathsf{P}^{\#\mathsf{P}}\subseteq\exists\mathsf{BPP}$, the equations of this variety do not have small, low degree constant free algebraic circuits. In a subsequent work ([BIL+19]), the results of [BIJL18] are generalized to min-rank or slice-rank varieties. However, in the bounded coefficient setting (and over finite fields), our results show that the contrary is true, and $\mathsf{VP}$ does have efficiently computable low degree equations. We also note that the set-up in these papers differ from that in our paper, and that of [GKSS17, FSV18]. One way to interpret this difference is that [BIJL18] shows that “variety of small completion rank tensors” cannot be “cut out” by efficient equations, whereas [GKSS17, FSV18] and our paper asks if every equation for this variety requires large complexity.
@@
For Boolean circuits, Chow [Cho11] shows a way of circumventing the natural proofs’ barrier in [RR97] by providing (under standard cryptographic assumptions) an explicit almost natural proof that is useful against $P/\operatorname{poly}$ as well as constructive in nearly linear time, but compromises on the largeness condition. Furthermore, Chow [Cho11] shows the unconditional existence of a natural property useful against $P/\operatorname{poly}$ (infinitely often) constructive in linear size that has a weakened largeness condition. In some sense, Theorem 1.7 and Theorem 1.6 are analogous to the work of Chow [Cho11], albeit in the algebraic world.
@@ 1.5 Proof ideas
In order to show that the equations are non-trivial, in the sense that there exist polynomials with bounded integer coefficients which do not pass this test, we need to show that there are nonzero polynomials with bounded integer coefficients which vanish everywhere on the hitting set $\pazocal{H}$. We show this via a well-known lemma of Siegel [^6] A statement of the lemma can be found here. Refer to [Sie14] for details., which uses a simple pigeon-hole argument to show that an under-determined system of homogeneous linear equations where the constraint matrix has small integer entries has a nonzero solution with small integer entries.
@@
As was observed in [FSV18, GKSS17], a lower bound for equations for a class of polynomials is equivalent to showing the existence of succinctly describable hitting sets for this class. For our proof of Theorem 1.10 we show that, assuming that the permanent is sufficiently hard, the coefficient vectors of polynomials in $\mathsf{VNP}$ form a hitting set for the class $\mathsf{VP}$. The connection between hardness and randomness in algebraic complexity is well known via a result of Kabanets and Impagliazzo [KI04], and we use this connection for our proof, along with some additional ideas. It is useful to note that the above-mentioned result of Kabanets and Impagliazzo [KI04] is essentially an algebraic analogue of the hardness vs randomness paradigm introduced by Nisan and Wigderson [NW94] in the boolean world. We briefly describe a high level sketch of our proof in a bit more detail now.
@@
For the proof of Theorem 1.10 (more formally, Theorem 5.8), we show that the points in the image of the map $\mathsf{Gen}_{\operatorname{Perm}}$, can be viewed as the coefficient vectors of polynomials in $\mathsf{VNP}$, or, equivalently in the terminology in [FSV18, GKSS17], that the Kabanets-Impagliazzo hitting set generator is $\mathsf{VNP}$-succinct. To this end, we work with a specific instantiation of the construction of the Kabanets-Impagliazzo generator where the underlying construction of combinatorial designs is based on Reed-Solomon codes. Although this is perhaps the most well known construction of combinatorial designs, there are other (and in some parameters, better) constructions known. However, our proof relies on the properties of this particular construction to obtain the succinct description. Our final proof is fairly short and elementary, and is based on extremely simple algebraic ideas and making generous use of the fact that we are trying to prove a lower bound for equations for $\mathsf{VNP}$ and not $\mathsf{VP}$.
For Theorem 1.13, we prove an approximative version of Theorem 5.8 by using the result of Bürgisser [Bür18] [^7] The original work is from 2004, this version corrects an error in one of the proofs. which showed that approximative circuits are particularly useful in computing factors of polynomials that have small size, irrespective of their degree.
@@
We begin with some notations and preliminaries in Section 2. In Section 3, we outline the existence of efficient hitting sets, over characteristic-zero fields, for any class that has a low-degree, low-variate universal polynomials; $\mathsf{VP}$ and $\mathsf{VNP}$ are examples of such classes. Existence of efficient hitting sets for $\mathsf{VP}$ and $\mathsf{VNP}$ over finite fields is outlined in Section 3.2. Note that all the arguments in Section 3 are present in previous works, but the results over characteristic zero have not been stated in this generality prior to this work, to the best of our knowledge. In Section 4, we use the above results to first prove Theorems 1.6 to 1.9. We then show the conditional hardness of equations for $\mathsf{VNP}$ in Section 5. Finally, we provide some open questions that arise from our work, and the prior literature around algebraic natural proofs in Section 6.
@@ 2.1 Notation and basics
Lemma 2.3 (Polynomial Identity Lemma [Ore22, DL78, Sch80, Zip79]).
@@ 2.2.2 Equations, natural proofs and succinct hitting sets over characteristic zero
$\mathsf{VP}$ has $\mathsf{VP}$-natural proofs $\Longleftrightarrow$ $\mathsf{VP}$ does not have $\mathsf{VP}$-succinct hitting sets.
@@ 2.3 Universal Circuits
Let $f$ be an $n$-variate degree $d$ polynomial computable by a circuit $C$ of size $s$. Using the classical depth reduction result due to Valiant et al. [VSBR83], $f$ has a circuit $C^{\prime}$ of size $s^{\prime}=\operatorname{poly}(n,d,s)$ and depth $\ell=O(\log d)$ with the following properties (see, e.g., [Sap15] for a complete proof).
@@ 2.4 Hardness-Randomness Connections
Kabanets and Impagliazzo [KI04] obtain hitting set generators from polynomials that are hard to compute for algebraic circuits. The following lemma is crucial to the proof of Theorem 5.8.
@@ 3.2 Finite fields: Hitting Sets for VP and VNP
Lemma 3.8 (Folklore (cf. Forbes [For14, Lemma 3.2.14])).
@@ 3.3 Finite fields: Hitting Sets for VNP
In order to prove the existence of a hitting set for the class $\pazocal{D}(n,d,s)$, we will need a bound on the number of polynomials in the class $\pazocal{D}(n,d,s)$ as well as a bound on the size of an explicit hitting set for the class of $n$-variate degree at most $d$ polynomials. These two bounds are summarized in the following claims, proofs of which can be found in [For14].
Claim 3.10 (Lemma 3.1.6 in [For14]).
@@
The rest of the proof follows exactly along the lines of the proof of Lemma 3.2.14 in [For14].